# SuperCarver: Texture-Consistent 3D Geometry Super-Resolution for High-Fidelity Surface Detail Generation

### v2

@@ IV-B Ablation Study
As reported in Table II, directly fine-tuning a standard diffusion model conditioned on color images lead to significantly degraded performance, with the MAE metric of 21.64. More importantly, as illustrated in Figure 9, the stochastic nature inevitably results in unstable predictions. As reported in the last two columns of Table II, the joint injection of semantic and geometric priors are indeed beneficial.
For inverse normal rendering, we experimented with another two distance field deformation schemes, i.e., fixing grids while optimizing distances; jointly updating both grids and distances. To obtain quantitative results, we computed the MAE metric of normal maps rendered from optimized meshes (note that Table I and Table II are targeted at diffused normal maps). As evidenced in Table III, the two variants show sub-optimal performances, in that only optimizing distances usually causes weakened details while the joint optimization strategy easily suffers from surface distortion especially when the diffused normal maps contain relatively larger errors.
Figure 9: Illustration of prediction variation issues of the stochastic baseline diffusion scheme as displayed in (d). The color image, ground-truth normal map, and the output of our original design are presented in (a), (b), and (c).
@@
Additionally, we quantitatively measured normal deviations across views in Table IV, in which the average MAE is 7.24. Comparing this metric with the overall normal prediction error with respect to ground-truth normal maps and the overall performances of other state-of-the-art competitors as reported in Table I, we can observe that our view consistency performance is satisfactory enough, as visualized in Figure 10.
IV-C Limitations & Future Works
@@
As the first learning-based solution for texture-consistent 3D geometry super-resolution, SuperCarver demonstrates satisfactory fidelity and applicability. Nevertheless, as a preliminary attempt in this direction, our current technical implementation still has several aspects of limitations for future explorations.
With finite image resolution, our normal boosting component cannot effectively enhance the detail patterns of elongated objects (e.g., knives, swords), since the rendered areas are too small to provide adequate appearance information. Localized processing would address the resolution issue.
Hindered by the inconsistent quality of training data, over-detailization occurs in certain cases. Theoretically, identifying geometry-irrelevant textures is supposed to be unified within the overall learning process when training with large volumes of data. Therefore, more stringent data selection protocols are needed to ensure geometry-texture consistency in the training samples. Still, it is worth noting that texture overfitting remains unresolved in the current 3D generation community. As shown in Figure 11, even the latest state-of-the-art foundation model Hunyuan3D [97] still suffers from the occurrence of overly detailed 3D geometric structures. (Note that SuperCarver works consistently well on Hunyuan3D-generated assets as illustrated in Figure 12.)
Additionally, the inherent properties of the adopted signed distance field representation restrict our capabilities to handle open surfaces and highly complicated topologies. Hence, we need to explore differentiable iso-surface extractors built upon more generic implicit geometry representations.
Notably, our second stage of mesh detail generation relies on iterative gradient-based optimization. One promising direction is to build feed-forward neural architectures to further improve the geometric quality and running efficiency.

### v3

@@ IV-B Ablation Study
As reported in Table III, directly fine-tuning a standard diffusion model conditioned on color images lead to significantly degraded performance, with the MAE metric of 21.64. More importantly, as illustrated in Figure 12, the stochastic nature inevitably results in unstable predictions. As reported in the last two columns of Table III, the joint injection of semantic and geometric priors are indeed beneficial.
For inverse normal rendering, we experimented with another two distance field deformation alternatives, i.e., fixing grids while optimizing distances; jointly updating both grids and distances. To produce quantitative results, we computed the MAE metric of normal maps rendered from optimized meshes (note that Table II and Table III are targeted at diffused normal maps). As evidenced in Table IV, the two variants lead to sub-optimal performances, in that only optimizing distances usually causes weakened details while the joint optimization strategy easily suffers from surface distortion especially when the generated normal maps contain relatively larger errors. Besides, although our targeted geometry super-resolution task devotes to generating realistic geometric details that represent mild deviation from the original surface, we still experimented with large deformation scenarios to explore the characteristics of our “value-fixing, grid-moving” strategy. As shown in Figure 13, our scheme cannot robustly deal with significant mesh deformations, i.e., producing obvious surface degradations and outliers. Through empirical explorations, our proposed strategy maintains satisfactory performance for vertex offsets bounded by approximately 0.03.
Additionally, we quantitatively measured normal deviations across views in Table V, in which the average MAE is 7.24. Comparing this metric with the overall normal prediction error with respect to ground-truth normal maps and the overall performances of other state-of-the-art competitors as reported in Table II, we can observe that our view consistency performance is satisfactory enough, as visualized in Figure 14.
IV-C Limitations, Failure Cases, and Future Works
@@
Figure 15: Results of applying SuperCarver on elongated objects.
Figure 16: Illustration of potential texture overfitting caused by SuperCarver. The first row shows three failure cases with obvious texture overfitting. The second row shows two examples where texture overfitting issues are basically negligible.
As a novel learning-based solution for texture-consistent 3D geometry super-resolution, SuperCarver demonstrates satisfactory fidelity and applicability. Nevertheless, as a preliminary attempt in this direction, our current technical implementation still has several aspects of limitations for future explorations.
With finite image resolution, our normal boosting component cannot sufficiently enhance the detail patterns of highly elongated objects (e.g., knives, swords), as shown in Figure 15. The reason is that the rendered areas are too small to provide adequate appearance information and visual cues. To overcome this issue, we might need to construct a localized processing pipeline where the original 3D mesh is rendered and detailized through adaptive sliding windows.
Hindered by inconsistent data quality of the training meshes, there inevitably exist over-detailization phenomena. As shown in Figure 16, our results show additional geometric details in the three cases of the first row. However, since our approach can implicitly learn discrimination capability from the training data distribution, it succeeds in avoiding texture overfitting in the last two cases of the second row. Theoretically, identifying geometry-irrelevant texture patterns is supposed to be unified within the overall learning process when training with large volumes of mesh data. Therefore, more stringent data selection protocols are required to ensure geometry-texture consistency in the training samples. In addition, in certain cases, the same texture semantically allows for multiple valid interpretations of 3D surface details. Although our deterministic framework significantly improves stability and multi-view inconsistency, it inherently sacrifices diversity to some extent when dealing with geometry-appearance ambiguity.
Nevertheless, it is worth mentioning that texture overfitting remains unresolved in the current 3D generation community. Even the current state-of-the-art proprietary foundation model [108] suffers from the occurrence of the overly detailized 3D geometric structures as depicted in Figure 17. (Note that our SuperCarver can still supplement higher level of surface details onto the output of such a powerful foundation model as shown in Figure 18.
Additionally, the inherent properties of our adopted signed distance field representation structure restrict our capabilities to handle open surfaces and/or highly complicated topologies. Hence, we need to explore differentiable iso-surface extractors built upon more generic implicit geometry representations such as appropriate adaptations of unsigned distance fields.
Finally, our second stage of mesh detail generation relies on iterative gradient-based optimization. One promising direction is to build feed-forward neural architectures to further improve the geometric quality and inference efficiency. Considering the normal detailization stage as a 2D branch and the subsequent mesh optimization stage as a 3D branch, we could attempt to re-design the 3D branch into a feed-forward architecture. For example, we could begin by extracting multi-view 2D features from both the rendered RGB images and the boosted normal maps, then back-projecting and merging pixel features to the 3D domain, forming a 3D feature volume. To achieve surface detailization, we could append a prediction head that consumes the grid features and directly regress the desired 3D grid offsets under the supervision of pre-computed ground-truth values.